As I clean my bedroom, I apply a constant force of 100 N with a speed of 2 m/s. What is my power rating?



1
Expert's answer
2021-02-22T10:23:58-0500

By the definition of the power, we have:


"P=\dfrac{W}{t}=\dfrac{Fd}{t}=Fv=100\ N\cdot2\ \dfrac{m}{s}=200\ W."